OpenStudy (anonymous):

If g(x)=2x + 1, find g(a + h)- g(a).

OpenStudy (anonymous):

g(a+h) = 2(a+h) + 1 = 2a+2h+1 g(a) = 2a +1 so: g(a+h)-g(a) = 2a+2h+1-(2a+1) = 2h